# Tensorization of neural networks for improved privacy and interpretability

### v2

@@ Abstract.
We present a tensorization algorithm for constructing tensor train representations of functions, drawing on sketching and cross interpolation ideas. The method only requires black-box access to the target function and a small set of sample points defining the domain of interest. Thus, it is particularly well-suited for machine learning models, where the domain of interest is naturally defined by the training dataset. We show that this approach can be used to enhance the privacy and interpretability of neural network models. Specifically, we apply our decomposition to (i) obfuscate neural networks whose parameters encode patterns tied to the training data distribution, and (ii) estimate topological phases of matter that are easily accessible from the tensor train representation. Additionally, we show that this tensorization can serve as an efficient initialization method for optimizing tensor trains in general settings, and that, for model compression, our algorithm achieves a superior trade-off between memory and time complexity compared to conventional tensorization methods of neural networks.
@@ 1. Introduction
Recently, TNs have been proposed as machine learning models for broader tasks. Initial studies in this area [40, 41] implemented 1D models, known as Tensor Trains (TTs) [42] or Matrix Product States (MPS) in the physics literature [43]. Subsequent research introduced the use of more complex tensor networks, such as Trees [44, 45], MERAs [46] or PEPS [47], for both supervised and unsupervised learning tasks.
@@ 1.2. Organization
In Section 2, we introduce the proposed tensorization algorithm, TT-RSS, with a step-by-step explanation provided in Section 2.2. Its continuous adaptation is discussed in Section 2.3, while its computational complexity is analyzed in Section 2.4. Additionally, we discuss the close relation between the sketching approach utilized and the Tensor Train Cross Interpolation algorithm [69] in Section 2.5.
@@ 1.3.1. Cross interpolation
Unlike the SVD, which identifies a set of vectors spanning the column and row spaces inherently, cross interpolation [70], also known as pseudoskeleton [71] or CUR decomposition [72], reconstructs a matrix using $r$ selected columns and $r$ selected rows. For a rank-$r$ matrix, this decomposition can be exact; for matrices of higher rank, the error is generally larger compared to SVD, but the method is significantly faster. Building on this idea, it was shown that iterative application of cross interpolation can decompose a high-dimensional tensor into a TT format [73]. To minimize the decomposition error, an additional optimization step is required to identify the $r$ columns and $r$ rows whose intersection forms the submatrix with maximum volume, a criterion proven to yield the lowest error [74]. As this combinatorial optimization is NP-hard, greedy algorithms have been developed to accelerate the process while maintaining practical performance [75]. This approach is commonly referred to as Tensor Train Cross Interpolation (TT-CI).
@@ 1.3.2. Sketching
In subsequent work, Hur et al. [63] applied a similar approach to decompose an empirical distribution $\hat{p}_{\textsc{e}}$, aiming to derive a TT approximation of the underlying density $p$. For efficiency, their algorithm assumes that the number of samples $N$ used to construct the empirical distribution is not excessively large, as the computation of random projections scales linearly with $N$. However, achieving a good approximation of $p$ requires $N$ to grow exponentially with the dimension $n$, making this method impractical for high-dimensional scenarios. The algorithm, Tensor Train via Recursive Sketching (TT-RS), has the notable advantage of being applicable to continuous functions, avoiding the need for explicit quantization by utilizing continuous embeddings.
@@ 1.4.2. Interpretability
Further work showed that injective, TI TTs [43] belong to the same phase if they can be interpolated without topological obstructions. However, when symmetry is imposed, the interpolation path is constrained and topological obstructions might occur, leading to phase transitions at the points where the interpolating TT becomes non-injective [103, 104]. These phases, known as Symmetry-Protected Topological (SPT) phases, are classified by the first and second cohomology groups for gapped quantum spin systems [106, 107].
@@ 1.5. Notations
Typically, functions are restricted to a subset defined by a collection of selected points, referred to as sketch samples or, more commonly, pivots, denoted as $\mathbf{x}=\{(\mathbf{x}_{1}^{i},\ldots,\mathbf{x}_{n}^{i}):\mathbf{x}_{j}^{i}\in X_{j}\}_{i=1}^{N}$. The set of possible values for the $k$-th variable in $\mathbf{x}$ is represented as $\mathbf{x}_{k}=\{\mathbf{x}_{k}^{i}\}_{i=1}^{N}$. For integers $a<b$, the notation $a\!:\!b$ represents the set $\{a,a+1,\ldots,b\}$, which is useful for describing subsets of variables. For example, $\mathbf{x}_{a:b}=\{(\mathbf{x}_{a}^{i},\ldots,\mathbf{x}_{b}^{i})\}_{i=1}^{N}$. We use $\mathbf{x}^{i}=(\mathbf{x}_{1}^{i},\ldots,\mathbf{x}_{n}^{i})$ to select a single point, and subscript notation may also be applied to specify subsets of variables.
@@ 2. Description of the algorithm
In this section we provide a detailed description of our proposed algorithm, Tensor Train via Recursive Sketching from Samples, or TT-RSS, presented in a similar form to the presentation of TT-RS in Ref. [63]. We first consider the simplified case of a discrete function $f:[d_{1}]\times\cdots\times[d_{n}]\to\mathbb{R}$ and assume it admits a TT representation with ranks $r_{1},\ldots,r_{n-1}$. Additionally, we are given a set of $N$ pivots, $\mathbf{x}=\{(\mathbf{x}_{1}^{i},\ldots,\mathbf{x}_{n}^{i}):\mathbf{x}_{j}^{i}\in[d_{j}]\}_{i=1}^{N}$, which, as stated previously, could represent training points from the model or points defining a region of interest within the domain of the function $f$.
@@ 2.1. Main idea
where $\alpha_{k}\in[r_{k}]$, and $\Phi_{k}$ and $\Psi_{k}$ are matrices whose column and row spaces, respectively, span the column and row spaces of $f$. Since we assume that $f$ admits a TT representation, one possible decomposition of its $k$-th unfolding matrix is obtained by contracting the first $k$ and the last $n-k$ cores, respectively:
(2.2) $\Phi_{k}(x_{1:k},\alpha_{k})$ $:=G(x_{1},\alpha_{1})G(\alpha_{1},x_{2},\alpha_{2})\cdots G(\alpha_{k-1},x_{k},\alpha_{k}),$
$\Psi_{k}(\alpha_{k},x_{k+1:n})$ $:=G(\alpha_{k},x_{k+1},\alpha_{k+1})\cdots G(\alpha_{n-2},x_{n-1},\alpha_{n-1})G(\alpha_{n-1},x_{n}),$
@@
(2.3) $\Phi_{k}(x_{1:k},\alpha_{k})=\Phi_{k}(x_{1:k-1},\alpha_{k-1})G(\alpha_{k-1},x_{k},\alpha_{k}),$
@@
When defining the CDEs (2.4), one possible approach would be to use the first $k-1$ cores obtained on the left-hand side to solve for the $k$-th core. However, introducing the matrices $\Phi_{k-1}$ and $\Phi_{k}$ on both sides of the equation helps mitigate the accumulation of errors that could arise from reusing the previously computed solutions $G_{1},\ldots,G_{k-1}$.
@@
  $s_{1},\ldots,s_{n-1},T_{2},\ldots,T_{n}\leftarrow\textsc{SketchForming}(\mathbf{x})$.
@@ 2.2. Details of subroutines
For clarity in the description of the entire algorithm, we divide it into the following subroutines: SketchForming (Section 2.2.1), Sketching (Section 2.2.2), Trimming (Section 2.2.3), SystemForming (Section 2.2.4), and Solving (Section 2.2.5).
2.2.1. SketchForming
@@
for $i\in[N_{a:b}]$, where $\mathbb{I}_{a:b}\in\mathbb{R}^{d_{a}\cdots d_{b}}$ denotes the identity matrix. This is, $P_{a:b}$ represents the orthogonal projection onto the subspace determined by the variables $x_{a}$ to $x_{b}$ of the pivots, $\mathbf{x}_{a:b}$. Similarly, we denote $P_{a:b}^{\intercal}:[d_{a}]\times\cdots\times[d_{b}]\times[N_{a:b}]\to\mathbb{R}$ as the function whose matricization is the transpose of the matricization of $P_{a:b}$, such that:
(2.9) $P_{a:b}^{\intercal}(x_{a:b},i)=P_{a:b}(i,x_{a:b}).$
@@
(2.10) $S_{k}(\beta_{k},x_{1:k})$ $:=P_{1:k}(\beta_{k},x_{1:k})=\mathbb{I}_{1:k}(\mathbf{x}^{\beta_{k}}_{1:k},x_{1:k}),$
$T_{k}(x_{k:n},\gamma_{k-1})$ $:=P_{k:n}^{\intercal}(x_{k:n},\eta_{k-1})U_{k}(\eta_{k-1},\gamma_{k-1})=\mathbb{I}_{k:n}(x_{k:n},\mathbf{x}^{\eta_{k-1}}_{k:n})U_{k}(\eta_{k-1},\gamma_{k-1}),$
where $U_{k}\in O(N_{k:n})$ is a random orthogonal (or, more generally, unitary) matrix sampled from the Haar measure as described in Ref. [109]. Note that with this definition, the sketches $S_{k}$ admit a recursive formulation if we define
@@
Choosing $U_{k}$ as an orthogonal random matrix aims to construct the sketches $T_{k}$ as Johnson-Lindenstrauss embeddings. The orthogonal projections $P^{\intercal}_{a:b}$ are deterministic matrices that restrict the action of $f$ to the subspace defined by $\mathbf{x}_{a:b}$. The isometry $U_{k}$ introduces randomness while preserving distances. No additional isometry is introduced in the left sketches $S_{k}$, as they are solely intended to reduce the number of equations, and introducing an isometry would be irrelevant for solving them.
@@ 2.2.2. Sketching
this is, evaluating the function $f$ on all the points in the set $\mathbf{x}_{1:k-1}\times[d_{k}]\times\mathbf{x}_{k+1:n}$. From $\bar{\Phi}_{k}$, we form the tensors $\widetilde{\Phi}_{k}$ by multiplying with $U_{k+1}$ on the right:
(2.13) $\widetilde{\Phi}_{k}(\beta_{k-1},x_{k},\gamma_{k}):=$ $\,\bar{\Phi}_{k}(\beta_{k-1},x_{k},\eta_{k})U_{k+1}(\eta_{k},\gamma_{k})$
@@
The tensors $\widetilde{\Phi}_{k}$ serve the same role as the tensors $\Phi_{k}$ on the right-hand side of Eq. 2.4, with the distinction that the indices $\beta_{k-1}$ and $\gamma_{k}$ range over $[N_{1:k-1}]$ and $[N_{k+1:n}]$, respectively, instead of $[r_{k-1}]$ and $[r_{k}]$. To form the tensors on the left-hand side of Eq. 2.4, one would typically contract $f$ with $S_{k-1}$ and $T_{k}$. However, as will be detailed in the subsequent subroutines, this additional computation can be avoided by reusing the previously computed $\widetilde{\Phi}_{k-1}$ and taking advantage of the recursive definition of $S_{k-1}$.
@@
In the general description of TT-RS provided in Ref. [63], left and right sketches are described as random projections, and only the construction of efficient sketches for Markov chains is discussed. However, applying random projections in a space of size $\prod_{i=1}^{n}d_{i}$ is generally infeasible. Our adaptation, TT-RSS, addresses this challenge by introducing the projections $P_{a:b}$, whose practical computation only requires evaluating $f$ on the selected pivots. This approach reduces the complexity of creating tensors $\bar{\Phi}_{k}$ to $O(N_{1:k-1}d_{k}N_{k+1:n})$.

### v3

@@ Abstract.
We present a tensorization algorithm for constructing tensor train/matrix product state (MPS) representations of functions, drawing on sketching and cross interpolation ideas. The method only requires black-box access to the target function and a small set of sample points defining the domain of interest. Thus, it is particularly well-suited for machine learning models, where the domain of interest is naturally defined by the training dataset. We show that this approach can be used to enhance the privacy and interpretability of neural network models. Specifically, we apply our decomposition to (i) obfuscate neural networks whose parameters encode patterns tied to the training data distribution, and (ii) estimate topological phases of matter that are easily accessible from the MPS representation. Additionally, we show that this tensorization can serve as an efficient initialization method for optimizing MPS in general settings, and that, for model compression, our algorithm achieves a superior trade-off between memory and time complexity compared to conventional tensorization methods of neural networks.
@@ 1. Introduction
Recently, TNs have been proposed as quantum-inspired machine learning models for broader tasks. Initial studies in this area [40, 41] implemented 1D models known as Matrix Product States (MPS) [42] or Tensor Trains (TTs) [43], the latter being the customary term in this context, which we will adopt throughout this work. Subsequent research introduced the use of more complex tensor networks, such as Trees [44, 45], MERAs [46] or PEPS [47], for both supervised and unsupervised learning tasks.
@@ 1.2. Organization
In Section 2, we introduce the proposed tensorization algorithm, TT-RSS, with a step-by-step explanation provided in Section 2.2. Its continuous adaptation is discussed in Section 2.3, while its computational complexity is analyzed in Section 2.4. Additionally, we discuss the close relation between the sketching approach utilized and the Tensor Train via Cross Interpolation algorithm [69] in Section 2.5.
@@ 1.3.1. Cross interpolation
Unlike the SVD, which identifies a set of vectors spanning the column and row spaces inherently, cross interpolation [70], also known as pseudoskeleton [71] or CUR decomposition [72], reconstructs a matrix using $r$ selected columns and $r$ selected rows. For a rank-$r$ matrix, this decomposition can be exact; for matrices of higher rank, the error is generally larger compared to SVD, but the method is significantly faster. Building on this idea, it was shown that iterative application of cross interpolation can decompose a high-dimensional tensor into a TT format [73]. To minimize the decomposition error, an additional optimization step—namely, the maxvol routine—is required to identify the $r$ columns and $r$ rows whose intersection forms the submatrix with maximum volume, a criterion proven to yield the lowest error [74]. As this combinatorial optimization is NP-hard, greedy algorithms have been developed to accelerate the process while maintaining practical performance [75]. This approach is commonly referred to as Tensor Train via Cross Interpolation (TT-CI).
@@ 1.3.2. Sketching
In subsequent work, Hur et al. [63] applied a similar approach to decompose an empirical distribution $\hat{p}_{\textsc{e}}$, aiming to derive a TT approximation of the underlying density $p$. For efficiency, their algorithm assumes that the number of samples $N$ used to construct the empirical distribution is not excessively large, as the computation of random projections scales linearly with $N$. However, achieving a good approximation of $p$ requires $N$ to grow exponentially with the dimension $n$, making this method impractical for high-dimensional scenarios. The algorithm, Tensor Train via Recursive Sketching (TT-RS), has the notable advantage of being natively applicable to continuous functions by utilizing continuous embeddings, thereby avoiding the need for explicit quantization.
@@ 1.4.2. Interpretability
Further work showed that injective, TI TTs [42] belong to the same phase if they can be interpolated without topological obstructions. However, when symmetry is imposed, the interpolation path is constrained and topological obstructions might occur, leading to phase transitions at the points where the interpolating TT becomes non-injective [103, 104]. These phases, known as Symmetry-Protected Topological (SPT) phases, are classified by the second cohomology group for gapped quantum spin systems [106, 107].
@@ 1.5. Notations
Typically, functions are restricted to a subset defined by a collection of selected points, referred to as sketch samples or, more commonly, pivots, denoted as $\mathbf{x}=\{(\mathbf{x}_{1}^{i},\ldots,\mathbf{x}_{n}^{i}):\mathbf{x}_{j}^{i}\in X_{j}\}_{i=1}^{N}$. The set of possible values for the $k$-th variable in $\mathbf{x}$ is represented as $\mathbf{x}_{k}=\{\mathbf{x}_{k}^{i}\}_{i=1}^{N}$. For integers $a<b$, the notation a:ba\!:\!b represents the set $\{a,a+1,\ldots,b\}$, which is useful for describing subsets of variables. For example, $\mathbf{x}_{a:b}=\{(\mathbf{x}_{a}^{i},\ldots,\mathbf{x}_{b}^{i})\}_{i=1}^{N}$. We use $\mathbf{x}^{i}=(\mathbf{x}_{1}^{i},\ldots,\mathbf{x}_{n}^{i})$ to select a single point, and subscript notation may also be applied to specify subsets of variables.
@@ 2. Description of the algorithm
In this section we provide a detailed description of our proposed algorithm, Tensor Train via Recursive Sketching from Samples, or TT-RSS, presented in a similar form to the presentation of TT-RS in Ref. [63]. We first consider the simplified case of a discrete function $f:[d_{1}]\times\cdots\times[d_{n}]\to\mathbb{R}$ and assume it admits a TT representation with ranks $r_{1},\ldots,r_{n-1}$. Additionally, we are given a set of NN pivots, $\mathbf{x}=\{(\mathbf{x}_{1}^{i},\ldots,\mathbf{x}_{n}^{i}):\mathbf{x}_{j}^{i}\in[d_{j}]\}_{i=1}^{N}$, which, as stated previously, could represent training points from the model or points defining a region of interest within the domain of the function $f$.
@@ 2.1. Main idea
where $\alpha_{k}\in[r_{k}]$, and $\Phi_{k}$ and $\Psi_{k}$ are matrices whose column and row spaces, respectively, span the column and row spaces of $f$. Since we assume that $f$ admits a TT representation, one possible decomposition of its $k$-th unfolding matrix is obtained by contracting the first $k$ and the last nkn-k cores, respectively:
(2.2) $\Phi_{k}(x_{1:k},\alpha_{k})$ $=G_{1}(x_{1},\alpha_{1})G_{2}(\alpha_{1},x_{2},\alpha_{2})\cdots G_{k}(\alpha_{k-1},x_{k},\alpha_{k}),$
$\Psi_{k}(\alpha_{k},x_{k+1:n})$ $=G_{k+1}(\alpha_{k},x_{k+1},\alpha_{k+1})\cdots G_{n-1}(\alpha_{n-2},x_{n-1},\alpha_{n-1})G_{n}(\alpha_{n-1},x_{n}),$
@@
(2.3) $\Phi_{k}(x_{1:k},\alpha_{k})=\Phi_{k-1}(x_{1:k-1},\alpha_{k-1})G_{k}(\alpha_{k-1},x_{k},\alpha_{k}),$
@@
When defining the CDEs (2.4), one possible approach would be to use the first k1k-1 cores obtained on the left-hand side to solve for the $k$-th core. However, introducing the matrices $\Phi_{k-1}$ and $\Phi_{k}$ on both sides of the equation helps mitigate the accumulation of errors that could arise from reusing the previously computed solutions $G_{1},\ldots,G_{k-1}$.
@@
In contrast to the TT-RS approach [63], which employs random sketches, we propose the use of orthogonal projection sketches. Specifically, we define sketches $S_{k-1}$ and $T_{k+1}$ such that applying them to $f$ corresponds to evaluating the function on the subsets of variables $\mathbf{x}_{1:k-1}$ and $\mathbf{x}_{k+1:n}$, respectively, thereby constructing the tensor $f(\mathbf{x}_{1:k-1},[d_{k}],\mathbf{x}_{k+1:n})$. This method is efficient for any function, requiring only a polynomial number of function calls, unlike random sketches, whose cost scales exponentially with $n$.
Although our method does not explore the full domain—as do random sketches in TT-RS or maxvol in TT-CI—we show that a small set of pivots from a relevant subregion suffices to achieve approximations at least as accurate as those from TT-CI, but at significantly lower cost (see Section 3). For our target applications, vanilla TT-RS is infeasible due to the exponential cost of random sketches. To select suitable pivots, we show that for regular functions, uniformly random points from the domain suffice, while for machine learning models—which are typically sparse tensors, or functions with flat landscapes and narrow peaks, often exhibiting discrete symmetries—using pivots from the training set is more effective. Prior work has also shown that in such settings, maxvol may be limited by ergodicity issues, and that incorporating domain knowledge is crucial to guide exploration toward relevant regions [77]. Our approach thus combines the strengths of TT-RS and TT-CI, overcoming the efficiency and exploration limitations of both.
@@
$s_{1},\ldots,s_{n-1},T_{2},\ldots,T_{n}\leftarrow\textsc{SketchBuilding}(\mathbf{x})$.
@@ 2.2. Details of subroutines
For clarity in the description of the entire algorithm, we divide it into the following subroutines: SketchBuilding (Section 2.2.1), Sketching (Section 2.2.2), Trimming (Section 2.2.3), SystemForming (Section 2.2.4), and Solving (Section 2.2.5).
2.2.1. SketchBuilding
@@
for $i\in[N_{a:b}]$, where $\mathbb{I}_{a:b}\in\mathbb{R}^{d_{a}\cdots d_{b}}$ denotes the identity matrix. This is, $P_{a:b}$ represents the orthogonal projection onto the subspace determined by the variables $x_{a}$ to $x_{b}$ of the pivots, $\mathbf{x}_{a:b}$. Similarly, we denote $P_{a:b}^{\top}:[d_{a}]\times\cdots\times[d_{b}]\times[N_{a:b}]\to\mathbb{R}$ as the function whose matricization is the transpose of the matricization of $P_{a:b}$, such that:
(2.9) $P_{a:b}^{\top}(x_{a:b},i)=P_{a:b}(i,x_{a:b}).$
The use of these projections to construct the sketches constitutes the main difference from the approach in Ref. [63], which relies solely on random sketching matrices.
@@
(2.10) $S_{k}(\beta_{k},x_{1:k})$ $=P_{1:k}(\beta_{k},x_{1:k})=\mathbb{I}_{1:k}(\mathbf{x}^{\beta_{k}}_{1:k},x_{1:k}),$
$T_{k}(x_{k:n},\gamma_{k-1})$ $=P_{k:n}^{\top}(x_{k:n},\eta_{k-1})U_{k}(\eta_{k-1},\gamma_{k-1})=\mathbb{I}_{k:n}(x_{k:n},\mathbf{x}^{\eta_{k-1}}_{k:n})U_{k}(\eta_{k-1},\gamma_{k-1}),$
where $U_{k}\in\mathrm{O}(N_{k:n})$ is a random orthogonal (or, in the case of complex coefficients, unitary) matrix sampled from the Haar measure as described in Ref. [109]. Note that with this definition, the sketches $S_{k}$ admit a recursive formulation if we define
@@
Choosing $U_{k}$ as an orthogonal random matrix aims to construct the sketches $T_{k}$ as Johnson-Lindenstrauss embeddings. Specifically, instead of applying the whole $U_{k}$, one can retain only a subset of its columns, yielding a projection to a lower-dimensional space, which helps reduce the computational cost of subsequent steps—such as the SVD in Trimming. Furthermore, since the orthogonal projections $P^{\top}_{a:b}$ are deterministic matrices that restrict the action of $f$ to the subspace defined by $\mathbf{x}_{a:b}$, the isometry $U_{k}$ introduces randomness while approximately preserving pairwise distances. No additional isometry is introduced in the left sketches $S_{k}$, as they are solely intended to reduce the number of equations, and introducing an isometry would be irrelevant for solving them.
@@ 2.2.2. Sketching
that is, evaluating the function $f$ on all the points in the set $\mathbf{x}_{1:k-1}\times[d_{k}]\times\mathbf{x}_{k+1:n}$, which entails a number of function calls that is polynomial in the number of pivots $N$ (see Section 2.4 for further details). The ability to substitute the projection with only a polynomial number of function calls is what makes our approach efficient, in contrast to Ref. [63], whose random sketches scale exponentially with $n$. From $\bar{\Phi}_{k}$, we form the tensors $\widetilde{\Phi}_{k}$ by multiplying with $U_{k+1}$ on the right:
(2.13) $\widetilde{\Phi}_{k}(\beta_{k-1},x_{k},\gamma_{k})$= $\,\bar{\Phi}_{k}(\beta_{k-1},x_{k},\eta_{k})U_{k+1}(\eta_{k},\gamma_{k})$
@@
The tensors $\widetilde{\Phi}_{k}$ serve the same role as the tensors $\Phi_{k}$ on the right-hand side of Eq. 2.4, with the distinction that the indices $\beta_{k-1}$ and $\gamma_{k}$ range over $N_{1:k-1}]$ and $N_{k+1:n}]$, respectively, instead of $r_{k-1}]$ and $r_{k}]$. To form the tensors on the left-hand side of Eq. 2.4, one would typically contract $f$ with $S_{k-1}$ and $T_{k}$. However, as will be detailed in the subsequent subroutines, this additional computation can be avoided by reusing the previously computed $\widetilde{\Phi}_{k-1}$ and taking advantage of the recursive definition of $S_{k-1}$.
@@
As commented earlier, in the general description of TT-RS provided in Ref. [63], left and right sketches are described as random projections, and only the construction of efficient sketches for Markov chains is discussed. However, applying random projections in a space of size $\prod_{i=1}^{n}d_{i}$ is generally infeasible. Our adaptation, TT-RSS, addresses this challenge by introducing the projections $P_{a:b}$, whose practical computation only requires evaluating $f$ on the selected pivots. This approach reduces the complexity of creating tensors $\bar{\Phi}_{k}$ to $O(N_{1:k-1}d_{k}N_{k+1:n})$.